OpenStudy (anonymous):

if 14/3 = x/y then 14/x=?

OpenStudy (anonymous):

14/3 = x/y (1/x)14/3 = (1/x)x/y 14/(3x) = 1/y 3(14/(3x)) = 3(1/y) 14/x = 3/y

OpenStudy (anonymous):

\[\frac{ 14 }{ 3 }=\frac{ x }{ y }=>14y=3x=>y=\frac{ 3x }{ 14 }=>\frac{ x }{ 14 }=\frac{ y }{ 3 }=>x =\frac{ 14y}{ 3 }=>\frac{ 1 }{ x }=\frac{ 14y }{ 3x^2 }=>\frac{ 14 }{ x }=\frac{ 14^2y }{ 3x^2 }\]

OpenStudy (anonymous):

was i wrong?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

o how

OpenStudy (anonymous):

nvm i got it

OpenStudy (anonymous):

Thank you so much!!!